# On the Curvature of Regge Metrics

### v1

@@ 2.5 Removing Frame-Dependency
This follows fairly simply from the definitions (21) and the change of basis formula (18). Let $\phi=\Psi_{f}^{-1}(\phi)$, which is a compatible test form. Then:
$\left\langle\hat{R}\wedge\hat{\phi}\right\rangle$ $=\left\langle\Psi_{f}^{-1}({f^{T}}^{*}\Omega)\wedge\Psi_{f}^{-1}(\hat{\phi})\right\rangle=\left\langle{f^{T}}^{*}\Omega\wedge\phi\right\rangle$
$\left\langle\hat{\mathrm{I\!I}}_{e}^{T}\wedge i_{e}^{*}\hat{\phi}^{T}\right\rangle$ $=\left\langle\Psi_{E_{e}^{T}}^{-1}(\hat{\mathrm{I\!I}}_{e}^{T})\wedge\Psi_{E_{e}^{T}}^{-1}(i_{e}^{*}\hat{\phi}^{T})\right\rangle=\left\langle\mathrm{I\!I}_{e}^{T}\wedge\mathrm{Ad}(\mu_{e}^{T})(i_{e}^{*}\phi)\right\rangle$
$\left\langle\hat{\Theta}_{p}^{T}\wedge i_{p}^{*}\hat{\phi}^{T}\right\rangle$ $=\left\langle\Psi_{E_{p,e}^{T}}^{-1}(\hat{\Theta}_{p}^{T})\wedge\Psi_{E_{p,e}^{T}}^{-1}(i_{p}^{*}\hat{\phi}^{T})\right\rangle=\left\langle\Theta_{p}w^{n-1}_{n}\wedge\mathrm{Ad}(A_{p,e}^{T}\mu_{e}^{T})(i_{p}^{*}\phi)\right\rangle.$
From here, we can apply the identity $\left\langle\alpha\wedge\mathrm{Ad}(h)(\beta)\right\rangle=\left\langle\mathrm{Ad}(h^{-1})(\alpha)\wedge\beta\right\rangle$ which is valid for $\mathfrak{gl}(n)$-valued forms. The only complication arises from the jump in fundamental form. Note that because $\mu_{e}^{T^{\prime}}=\begin{bmatrix}I&0\\ 0&-1\end{bmatrix}\mu_{e}^{T}$, we have
$\left\langle\mathrm{Ad}((\mu_{e}^{T})^{-1})(\mathrm{I\!I}_{e}^{T})\wedge i_{e}^{*}\phi\right\rangle$- $\left\langle\mathrm{Ad}((\mu_{e}^{T^{\prime}})^{-1})(\mathrm{I\!I}_{e}^{T^{\prime}})\wedge i_{e}^{*}\phi\right\rangle$
$=\left\langle\mathrm{Ad}((\mu_{e}^{T})^{-1})(\mathrm{I\!I}_{e}^{T})-\mathrm{Ad}((\mu_{e}^{T})^{-1})(-\mathrm{I\!I}_{e}^{T^{\prime}})\wedge i_{e}^{*}\phi\right\rangle$
$=\left\langle\mathrm{Ad}((\mu_{e}^{T})^{-1})([\![\mathrm{I\!I}_{e}]\!])\wedge i_{e}^{*}\phi\right\rangle$
Note that, since $i_{\mathring{p}}^{*}\hat{\phi}^{T}$ is not single-valued due to (20), at first glance the last term seems to have some ambiguity. However, per the preceding paragraph, $\hat{\Theta}_{p}^{T}$ is invariant under changes of the normal vectors $\vec{\nu}$ and $\vec{n}$. In other words, if $h=\exp\left(s~r_{p,e}^{T}(x)w^{n-1}_{n}\right)$, then $\hat{\Theta}_{p}^{T}]_{E_{p,e}^{T}}=[\hat{\Theta_{p}^{T}}]_{E_{p,e}^{T}\cdot h^{-1}}$. So, because ${\chi_{s,p}^{T}}^{*}\hat{\phi}^{T}]_{E_{p,e}^{T}}=\mathrm{Ad}(h)([{\chi_{0,p}^{T}}^{*}\hat{\phi}^{T}]_{E_{p,e}^{T}})$ by (20) and $\Psi_{E_{p,e}^{T}}$ is an isometry, we have
@@
So, in (22), $i_{\mathring{p}}^{*}\hat{\phi}^{T}$” could be thought of as a shorthand for ${\chi_{s,p}^{T}}^{*}\hat{\phi}^{T}$, where $s$ may be any convenient number in $[0,1]$. Concretely, this means $x\in\mathring{p}$ can be approached along a ray of any convenient angle, and the result will be the same. Additionally, due to equation (15), this term does not depend on the choice of $T$ of which $p$ is a face.
The right side of equation (22) does not depend on $f$ at all, so as long as $\hat{\phi}\in\mathcal{A}(f)$ for some compatible frame $f$, this definition can be used to compute the distributional curvature. The caveats on the integrals are the same as for (1): the integrals over $\mathring{e}$ do not depend on the choice of $T$, but care must be chosen so that the integral is evaluated with the orientation induced on $e$ by the orientation of $T$.
This expression has some close similarities with the densitized distributional curvature investigated in [8]. In fact, the two expressions are equivalent, with the main difference being the choice of how to represent the second fundamental form and which indices are raised/lowered. See proposition (1) in the appendix for a proof. As explained in [8], various traces of this distribution can be taken to obtain the Ricci curvature, Einstein tensor, and the scalar curvature.
@@ 3 Construction of Compatible Frames
Suppose there exists a $C^{2}$ homotopy of Regge metrics $g(t)$ such that $g(1)=g$ and $g(0)=g_{0}$ is a smooth metric, and there exists a smooth $g_{0}$-orthonormal frame $f_{0}$. Then there exists a $C^{2}$ homotopy of frames $f(t)$ such that $f(0)=f_{0}$, $f^{T}(t)$ is $g(t)$-orthonormal when restricted to $T$, and $f(t)$ satisfies conditions 1, 2, and 3 of compatible frames.
@@
First, we need a result that allows us to extend frames. This is a technical result that ultimately relies on the fact that functions can be extended smoothly from closed sets. We will specifically use the theorem as stated in [18] (see also [11, Lemma 2.26]), which can be simplified for our purposes in the way detailed below. In the remainder of this section, we will stop abusing terminology and use the word “polytope” to refer to a genuine polytope (as opposed to the image of a polytope under a smooth embedding). Our reason for doing so is that the extension procedure that we will soon describe takes place on the blow-up $B_{T}$, which we will assume to be a polytope. (In fact we will assume it to be convex.) We will use the notation $\hat{T}$ for polytopes below; ultimately we will take $\hat{T}=B_{T}$ when we begin constructing a compatible frame.
@@
In both cases, $\|x-x_{0}\|$ is bounded by a constant multiple of $\|\vec{u}\|$, which is in turn bounded by a constant multiple of $|\left\langle\vec{n}_{i},\vec{u}\right\rangle|+|\left\langle\vec{n}_{j},\vec{u}\right\rangle|$, since $(\vec{n}_{i},\vec{n}_{j})$ are linearly independent and normal to $\hat{p}_{ij}$. $\lambda^{i}(x)$ is exactly equal to $\left\langle\vec{n}_{i},\vec{u}\right\rangle$ since $x_{0}\in e_{i}$, and similarly for $\lambda^{j}$. Therefore there is a constant $C$ such that $\|x-x_{0}\|\leq C(\lambda^{i}(x)+\lambda^{j}(x))$.
@@
$=(u^{-1}\dot{u})^{m}_{i}\tilde{g}_{mj}+(u^{-1}\dot{u})^{m}_{j}\tilde{g}_{mi}+(u^{\intercal}\tilde{\sigma}(t)u)^{j}_{i}$
@@
In the next couple of paragraphs, we will argue that (23) admits a unique solution $u:[0,1]\times\hat{T}\to GL(n)$ for any Lipschitz choice of $K$. We start by constructing a solution for one particular choice of $K$, and then we leverage it to construct solutions for other choices of $K$.
@@
Let $T$, $\hat{T}$, $\Phi$, $g(t)$, and $f_{0}$ be as above. Then there exists a $C^{2}$ homotopy of $C^{2}$ frames $f(t):T\to\mathcal{F}_{GL}(T)$ such that $f(0)=f_{0}$ and $\frac{\partial}{\partial t}g(t)(f_{i}(t),f_{j}(t))=0$.
@@
In the proof of Lemma 3, $\hat{T}$ can be set equal to $T$ and $\Phi$ can simply be the identity map. Then choosing a $C^{2}$ map $K:[0,1]\times T\to\mathfrak{so}(n)$ is enough to produce a solvable o.d.e. for $u$. $K=0$ is a valid choice. ∎
@@
Now we bring it all together. $F^{T}(t)|_{p}$ and $F^{T}(t)|_{e}$ are defined and $C^{2}$ on faces of the polytope $B_{T}\subseteq\mathbb{R}^{n}$ and single-valued everywhere, therefore they can be extended, for each $t$, to a section $F^{T}(t):B_{T}\to{\Phi^{T}}^{*}\mathcal{F}_{O}(T)$ which remains orthonormal for all $t$, is $C^{2}$ on $B_{T}\backslash E_{B_{T}}$, and has bounded derivatives on $\mathring{B_{T}}$. This frame is $C^{2}$ in time and satisfies $F^{T}(0)={\Phi^{T}}^{*}f_{0}$. Lastly, set $f^{T}(t)(x)=F^{T}(t)(\Phi^{-1}(x))$ for $x\in S_{0}(T)\cup S_{1}(T)$. The resulting frame satisfies properties 1-3 of a $g(t)$-compatible frame by construction, and $f^{T}(0)=f_{0}$, so $f^{T}(1)$ is a compatible frame. ∎
@@ 4 Generalized Gauss-Bonnet Theorem
When specialized to 2 dimensions, the expression (1) simplifies considerably, since $\mathfrak{so}(2)$ and $\Lambda^{2}(M)$ are both one-dimensional. In this case there exists a function $K^{T}:T\to\mathbb{R}$ such that $\Omega^{T}=K^{T}w^{2}_{1}\otimes dA^{T}$, where $dA^{T}$ is the positively oriented volume form on $T$ induced by the metric $g_{T}$. $K^{T}$ is precisely the Gauss curvature, and does not depend on $f^{T}$. Additionally, the adjoint action is trivial on $\mathfrak{so}(2)$, and if $h$ is a smooth $SO(2)$-valued function which is a rotation by the angle $\theta$, then $h^{-1}dh$ is equal to $\begin{bmatrix}0&-d\theta\\ d\theta&0\end{bmatrix}=w^{1}_{2}\otimes d\theta$.
@@
This quantity is frame-independent and identical to the densitized distributional Gauss curvature investigated in [5, 2]. The fact that the frame dependence of the distributional curvature is concentrated in boundary terms is reminiscent of the following formula that is valid for a smooth metric and smooth frame, obtained from integration by parts:
$\int_{M}f^{*}\omega^{1}_{2}\wedge d\phi=\int_{M}\phi KdA-\int_{\partial M\backslash E_{M}}\phi f^{*}\omega^{1}_{2}.$
Therefore we can interpret the boundary components of the distributional curvature as being a distributional version of the connection one-form pulled back to the boundary. They read
$\sum_{e\subset\partial M}\int_{\mathring{e}}\phi(kds-d\bar{\mu}_{e})-\sum_{p\in\partial M}([\![\mu]\!]|_{p}+\pi-\sum_{T\ni p}\theta_{p}^{T})\phi(p),$
@@
Note that, in the case that $g$ is a smooth metric, this theorem reduces to the classical Gauss-Bonnet theorem. Also note that we are using slightly different sign conventions for the geodesic curvature $k$ than some authors do.
@@
$=\sum_{T\subseteq M}\int_{\mathring{T}}\left\langle{f^{T}}^{*}\omega\wedge\frac{1}{2}w^{2}_{1}\otimes d(1)\right\rangle+\sum_{e\subset\partial M}\int_{\mathring{e}}d\bar{\mu}_{e}-\sum_{p\in\partial M}\phi(p)[\![\mu]\!]|_{p}$
@@
Appendix A Appendix: Nice Manifolds, Blow-Ups, and Integration by Parts

### v2

@@ 2.5 Removing Frame-Dependency
[…]
$\left\langle\hat{R}\wedge\hat{\phi}\right\rangle$ $=\left\langle\Psi_{f}^{-1}(\hat{R})\wedge\Psi_{f}^{-1}(\hat{\phi})\right\rangle=\left\langle{f^{T}}^{*}\Omega\wedge\phi\right\rangle$
$\left\langle\hat{\mathrm{I\!I}}_{e}^{T}\wedge i_{\mathring{e}}^{*}\hat{\phi}|_{T}\right\rangle$ $=\left\langle\Psi_{E_{e}^{T}}^{-1}(\hat{\mathrm{I\!I}}_{e}^{T})\wedge\Psi_{E_{e}^{T}}^{-1}(i_{\mathring{e}}^{*}\hat{\phi}|_{T})\right\rangle=\left\langle\mathrm{I\!I}_{e}^{T}\wedge\mathrm{Ad}(\mu_{e}^{T})(i_{\mathring{e}}^{*}\phi)\right\rangle$
$\left\langle\hat{\Theta}_{p}^{T}\wedge i_{\mathring{p}}^{*}\hat{\phi}|_{T}\right\rangle$ $=\left\langle\Psi_{E_{p,e}^{T}}^{-1}(\hat{\Theta}_{p}^{T})\wedge\Psi_{E_{p,e}^{T}}^{-1}(i_{\mathring{p}}^{*}\hat{\phi}|_{T})\right\rangle=\left\langle\Theta_{p}w^{n-1}_{n}\wedge\mathrm{Ad}(A_{p,e}^{T}\mu_{e}^{T})(i_{\mathring{p}}^{*}\phi)\right\rangle.$
From here, we can apply the identity $\left\langle\alpha\wedge\mathrm{Ad}(h)(\beta)\right\rangle=\left\langle\mathrm{Ad}(h^{-1})(\alpha)\wedge\beta\right\rangle$ which is valid for $\mathfrak{so}(n)$-valued forms. The only complication arises from the jump in second fundamental form. Note that because $\mu_{e}^{T^{\prime}}=\begin{bmatrix}I&0\\ 0&-1\end{bmatrix}\mu_{e}^{T}$, we have
$\left\langle\mathrm{Ad}((\mu_{e}^{T})^{-1})(\mathrm{I\!I}_{e}^{T})\wedge i_{\mathring{e}}^{*}\phi\right\rangle$- $\left\langle\mathrm{Ad}((\mu_{e}^{T^{\prime}})^{-1})(\mathrm{I\!I}_{e}^{T^{\prime}})\wedge i_{\mathring{e}}^{*}\phi\right\rangle$
$=\left\langle\mathrm{Ad}((\mu_{e}^{T})^{-1})(\mathrm{I\!I}_{e}^{T})-\mathrm{Ad}((\mu_{e}^{T})^{-1})(-\mathrm{I\!I}_{e}^{T^{\prime}})\wedge i_{\mathring{e}}^{*}\phi\right\rangle$
$=\left\langle\mathrm{Ad}((\mu_{e}^{T})^{-1})([\![\mathrm{I\!I}_{e}]\!])\wedge i_{\mathring{e}}^{*}\phi\right\rangle$
Note that, since $i_{\mathring{p}}^{*}\hat{\phi}|_{T}$ is not single-valued due to (20), at first glance the last term seems to have some ambiguity. However, $\hat{\Theta}_{p}^{T}$ is invariant under rotations of the normal vectors $\vec{\nu}$ and $\vec{n}$ with which it is defined. More precisely, if $h=\exp\left(s~r_{p,e}^{T}(x)w^{n-1}_{n}\right)$, then $\hat{\Theta}_{p}^{T}]_{E_{p,e}^{T}}=[\hat{\Theta}_{p}^{T}]_{E_{p,e}^{T}\cdot h^{-1}}$. So, because ${\chi_{s,p}^{T}}^{*}\hat{\phi}^{T}]_{E_{p,e}^{T}}=\mathrm{Ad}(h)([{\chi_{0,p}^{T}}^{*}\hat{\phi}^{T}]_{E_{p,e}^{T}})$ by (20) and $\Psi_{E_{p,e}^{T}}$ is an isometry, we have
@@
So, in (22), $i_{\mathring{p}}^{*}\hat{\phi}|_{T}$” could be thought of as a shorthand for ${\chi_{s,p}^{T}}^{*}\hat{\phi}^{T}$, where $s$ may be any convenient number in $[0,1]$. Concretely, this means each $x\in\mathring{p}$ can be approached along a ray of any convenient angle, and the result will be the same. Additionally, due to equation (15), this term does not depend on the choice of $T$ of which $p$ is a face.
The right side of equation (22) does not depend on $f$ at all, so as long as $\hat{\phi}\in\mathcal{A}(f,\mathcal{T},M)$ for some compatible frame $f$, this definition can be used to compute the distributional curvature. The caveats on the integrals are the same as for (4): the integrals over $\mathring{e}$ do not depend on the choice of $T$, but care must be chosen so that the integral is evaluated with the orientation induced on $e$ by the orientation of $T$, and the integrals over $\mathring{p}$ must be evaluated using the orientation induced from the face $e$ such that $E_{p,e}^{T}$ has an inward-pointing normal vector as its last entry.
This expression has some close similarities with the densitized distributional curvature investigated in [13]. In fact, the two expressions are equivalent, with the main difference being the choice of how to represent the second fundamental form and which indices are raised/lowered. See Appendix B for a proof. As explained in [13], various traces of this distribution can be taken to obtain the Ricci curvature, Einstein tensor, and the scalar curvature.
@@ 3 Construction of Compatible Frames
Let $g$ be a Regge metric for the mesh $\mathcal{T}$ such that each polytope $T\subseteq M$ has a blow-up which is a closed convex polytope in $\mathbb{R}^{n}$, and suppose there exists a $C^{2}$ homotopy of Regge metrics $g(t)$ such that $g(1)=g$, $g(0)=:g_{0}$ is a smooth metric, and there exists a smooth $g_{0}$-orthonormal frame $f_{0}$. Then there exists a $C^{2}$ homotopy of frames $f(t)$ such that $f(0)=f_{0}$, $f^{T}(t)$ is $g(t)$-orthonormal when restricted to each $T$, $f(t)$ satisfies conditions 1, 2, and 3 of compatible frames, and the maps $(t,x)\mapsto F^{T}(t)(x)$ vary continuously as maps $0,1]\times B_{T}\to\mathcal{F}_{GL}(T)$, where $F^{T}$ is the blown-up frame from condition 1.
@@
First, we need a result that allows us to extend frames. This is a technical result that ultimately relies on the fact that functions can be extended smoothly from closed sets. We will specifically use the theorem as stated in [24] (see also [16, Lemma 2.26]), which can be simplified for our purposes in the way detailed below. In the remainder of this section, we will stop abusing terminology and use the word “polytope” to refer to a genuine polytope (as opposed to the image of a polytope under a smooth embedding). Our reason for doing so is that the extension procedure that we will soon describe takes place on the blow-up $B_{T}$, which we will assume to be a polytope. (In fact we will assume it to be convex.) We will use the notation $\hat{T}$ for polytopes below; ultimately we will take $\hat{T}=B_{T}$ when we begin constructing a compatible frame.
@@
In both cases, $\|x-x_{0}\|$ is bounded by a constant multiple of $\|\vec{u}\|$, which is in turn bounded by a constant multiple of $|\left\langle\vec{n}_{i},\vec{u}\right\rangle|+|\left\langle\vec{n}_{j},\vec{u}\right\rangle|$, since $(\vec{n}_{i},\vec{n}_{j})$ are linearly independent and normal to $T\hat{p}_{ij}$. $\lambda^{i}(x)$ is exactly equal to $\left\langle\vec{n}_{i},\vec{u}\right\rangle$ since $x_{0}\in e_{i}$, and similarly for $\lambda^{j}$. Therefore there is a constant $C$ such that $\|x-x_{0}\|\leq C(\lambda^{i}(x)+\lambda^{j}(x))$.
@@
  1. 1.
    There are $C^{2}$ (in both space and time) frames $f(t)|_{\hat{e}_{i}}:\hat{e}_{i}\to\Phi^{*}\mathcal{F}_{GL}(T)$, defined on some subset $\{\hat{e}_{i}\}_{i=1}^{N}$ of the codimension-1 faces of $\hat{T}$.
  2. 2.
    For each $i,j\in 1,\dots,N$, $f(t)|_{\hat{e}_{i}}(x)=f(t)|_{\hat{e}_{j}}(x)$ for all $t\in[0,1]$ and $x\in\hat{p}_{ij}=\hat{e}_{i}\cap\hat{e}_{j}$, and $\frac{d}{dt}[g(t)((f(t)|_{\hat{e}_{i}})_{j},(f(t)|_{\hat{e}_{i}})_{k})]=0$ for all $i,j,k,t$.
  3. 3.
    There is a smooth $g_{0}$-orthonormal frame $f_{0}:\hat{T}\to\Phi^{*}\mathcal{F}_{GL}(T)$ such that $f_{0}(x)=f(0)|_{\hat{e}_{i}}(x)$ for all $i$ and all $x\in\hat{e}_{i}$.
[…]
@@
$=(u^{-1}\dot{u})^{m}_{i}\tilde{g}_{mj}+(u^{-1}\dot{u})^{m}_{j}\tilde{g}_{mi}+(u^{\intercal}\tilde{\sigma}(t)u)_{ij}$
@@
In the next few paragraphs, we will argue that (23) admits a unique solution $u:[0,1]\times\hat{T}\to GL(n)$ for any Lipschitz choice of $K$. We start by constructing a solution for one particular choice of $K$, and then we leverage it to construct solutions for other choices of $K$.
@@
Let $T$, $g(t)$, and $f_{0}$ be as above. Then there exists a $C^{2}$ homotopy of $C^{2}$ frames $f(t):T\to\mathcal{F}_{GL}(T)$ such that $f(0)=f_{0}$ and $\frac{\partial}{\partial t}g(t)(f_{i}(t),f_{j}(t))=0$.
@@
In the proof of Lemma 13, $\hat{T}$ can be set equal to $T$ and $\Phi$ can simply be the identity map. We no longer have any boundary conditions on $f(t)|_{\hat{e}_{i}}$. Then choosing a $C^{2}$ map $K:[0,1]\times T\to\mathfrak{so}(n)$ is enough to produce a solvable o.d.e. for $u$. $K=0$ is a valid choice. ∎
@@
Now we bring it all together. $F^{T}(t)|_{p}$ and $F^{T}(t)|_{e}$ are defined and $C^{2}$ on faces of the polytope $B_{T}\subseteq\mathbb{R}^{n}$ and single-valued everywhere, therefore they can be extended, for each $t$, to a section $F^{T}(t):B_{T}\to{\Phi^{T}}^{*}\mathcal{F}_{O}(T)$ which remains orthonormal for all $t$, is $C^{2}$ on $B_{T}\backslash E_{B_{T}}$, and has bounded derivatives on $\mathring{B}_{T}$. This frame is $C^{2}$ in time and satisfies $F^{T}(0)={\Phi^{T}}^{*}f_{0}$. Lastly, set $f^{T}(t)(x)=F^{T}(t)(\Phi^{-1}(x))$ for $x\in S_{0}(T)\cup S_{1}(T)$. The resulting frame satisfies properties 1-3 of a $g(t)$-compatible frame by construction, and $f^{T}(0)=f_{0}$, so $f(1)$ is a compatible frame. ∎
@@ 4 Generalized Gauss-Bonnet Theorem
When specialized to 2 dimensions, the expression (4) simplifies considerably, since $\mathfrak{so}(2)$ and $\Lambda^{2}(M)$ are both one-dimensional. In this case there exists a function $K^{T}:T\to\mathbb{R}$ such that $\Omega^{T}=K^{T}w^{2}_{1}\otimes dA^{T}$, where $dA^{T}$ is the positively oriented volume form on $T$ induced by the metric $g^{T}$. $K^{T}$ is precisely the Gauss curvature, and does not depend on $f^{T}$. Additionally, the adjoint action is trivial on $\mathfrak{so}(2)$, and if $h$ is a smooth $SO(2)$-valued function which is a rotation by the angle $\theta$, then $h^{-1}dh$ is equal to $\begin{bmatrix}0&-d\theta\\ d\theta&0\end{bmatrix}=w^{1}_{2}\otimes d\theta$.
@@
This quantity is frame-independent and identical to the densitized distributional Gauss curvature investigated in [12, 9, 1]. The fact that the frame dependence of the distributional curvature is concentrated in boundary terms is reminiscent of the following formula that is valid for a smooth metric and smooth frame, obtained from integration by parts:
$\int_{M}f^{*}\omega^{1}_{2}\wedge d\phi=\int_{\mathring{M}}\phi KdA-\int_{\partial M\backslash E_{M}}\phi f^{*}\omega^{1}_{2}.$
Therefore we can interpret the boundary components of the distributional curvature as being a distributional version of the connection one-form pulled back to the boundary. The analogue of $\int_{\partial M\backslash E_{M}}\phi f^{*}\omega^{1}_{2}$ above would be given by
$\sum_{e\subset\partial M}\int_{\mathring{e}}\phi(d\bar{\mu}_{e}-kds)-\sum_{p\in\partial M}([\![\mu]\!]|_{p}+\pi-\sum_{T\ni p}\theta_{p}^{T})\phi(p),$
@@
Note that, in the case that $g$ is a smooth metric, this theorem reduces to the classical Gauss-Bonnet theorem. Also note that we are using different sign conventions for the geodesic curvature $k$ than some authors do.
@@
$=\sum_{T\subseteq M}\int_{\mathring{T}}\left\langle{f^{T}}^{*}\omega\wedge\frac{1}{2}w^{2}_{1}\otimes d(1)\right\rangle+\sum_{e\subset\partial M}\int_{\mathring{e}}d\bar{\mu}_{e}-\sum_{p\in\partial M}[\![\mu]\!]|_{p}$
@@
Appendix A Appendix: Polyhedral Manifolds, Blow-Ups, and Integration by Parts